# A Diffusion-based generative equalizer for music restoration

### v1

@@ Abstract
This paper presents a novel approach to audio restoration, focusing on the enhancement of low-quality music recordings, and in particular historical ones. Building upon a previous algorithm called BABE, or Blind Audio Bandwidth Extension, we introduce BABE-2, which presents a series of significant improvements. This research broadens the concept of bandwidth extension to generative equalization, a novel task that, to the best of our knowledge, has not been explicitly addressed in previous studies. BABE-2 is built around an optimization algorithm utilizing priors from diffusion models, which are trained or fine-tuned using a curated set of high-quality music tracks. The algorithm simultaneously performs two critical tasks: estimation of the filter degradation magnitude response and hallucination of the restored audio. The proposed method is objectively evaluated on historical piano recordings, showing a marked enhancement over the prior version. The method yields similarly impressive results in rejuvenating the works of renowned vocalists Enrico Caruso and Nellie Melba. This research represents an advancement in the practical restoration of historical music. Historical music restoration examples are available at: research.spa.aalto.fi/publications/papers/dafx-babe2/.
@@ 1 Introduction
Historical music recordings suffer from severe impairment due to limitations of the physical recording media, alongside the wear and tear from playback and storage [1, 2]. These degradations can take the form of additive disturbances, decreased bandwidth, coloration, and nonlinear effects [3]. The ultimate goal is to enhance the quality of historical music recordings as to meet the audio standards of today, making it more accessible for people used to modern sound quality. Restoring these recordings is challenging for several reasons. First, the problem is ill-posed, meaning there are many potential solutions, which complicates the search for the correct one. Additionally, the degradation process is unknown, making this a blind inverse problem for which the solution must be found without definite information on the necessary repairs.
In the 1970s, Stockham et al. began groundbreaking efforts in digitally restoring gramophone recordings, notably those of the celebrated singer Enrico Caruso [4]. At the time, the approach was limited by technological constraints, relying on the estimation of an equalization curve derived from the spectral average of a contemporary recording. Since then, the field of audio restoration has evolved significantly [1]. Today, data-driven techniques employing deep neural networks can remove all additive disturbances simultaneously [5]. Additionally, deep generative models, such as diffusion models [6], address complex restoration tasks, such as bandwidth extension [7], with unprecedented success. This study investigates a critical question: “Can deep generative models elevate the quality of historical music recordings to modern standards?
To aid in this investigation, we introduce and explore the concept of generative equalization, a task designed to overcome the limitations of traditional equalization. Traditional equalizers face limitations on ill-conditioned scenarios, i.e., when some frequency bands are not present in the original recording. In such cases, the equalization curve must be cropped, otherwise, it will only amplify background noise. In contrast, a generative equalizer can synthesize the missing spectral components as to achieve the target spectral profile, as it is a generalization of bandwidth extension.
@@
The remaining of this paper is structured as follows. Sec. 2 presents sufficient information to follow the description of the new algorithm. There, the basics of diffusion models, diffusion posterior sampling, and the previous BABE algorithm are described. Following this, Sec. 3 delves into the enhancements made in the BABE-2 method, expanding the bandwidth extension techniques introduced in [7] to encompass the broader concept of generative equalization. Sec. 4 reports on experiments conducted with historical recordings of piano and singing voice. The discussion in Sec. 5 describes our procedure for selecting the training data used for restoring specific historical singers, and elucidates on some of the capabilities and limitations of BABE-2 for singing voice restoration. Finally, Sec. 6 concludes the paper.
@@ 2.1 Diffusion Models
$d\mathbf{x}=-\tau\nabla_{\mathbf{x}_{\tau}}\log p_{\tau}(\mathbf{x}_{\tau})d\tau.$ (1)
@@
$\mathbb{E}_{\mathbf{x}_{0}\sim p_{\text{data}},\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\lambda(\tau)\lVert D_{\theta}(\mathbf{x}_{0}+\tau\mathbf{\epsilon},\tau)-\mathbf{x}_{0}\rVert_{2}^{2}\right].$ (3)
In the subsequent sections, the term $\hat{\mathbf{x}}_{0}$ denotes the denoised estimate obtained through the application of neural network $D_{\theta}$.
@@ 2.2 Diffusion Posterior Sampling
where $\xi(\tau)$ serves as a gradient normalization factor [10].
Within this framework, $\hat{\mathbf{x}}_{0}=D_{\theta}(\mathbf{x}_{\tau},\tau)$ represents the Minimum Mean Squared Error (MMSE) estimate of the clean audio $\mathbf{x}_{0}$ given the noisy or transformed state $\mathbf{x}_{\tau}$. It is important to note that implementing this method necessitates differentiating through the denoiser, which is a deep neural network, and adds a layer of complexity to the process. Among other applications, this approach has been successfully used for solving audio restoration tasks, such as bandwidth extension, inpainting, and declipping [10].
@@ 2.3 Application to Blind Inverse Problems
In BABE, the filter design is constrained to yield a lowpass filter by projecting the parameters to ensure they describe a strictly decaying function [7]. This choice stabilizes the optimization process, as the primary focus is on responses attenuating the high frequencies. Such a parameterization, however, does not accommodate spectral coloration or resonances, which are common in historical recordings and can significantly impact the audio quality. We hypothesize that the constrained modeling capability degrades the performance of posterior sampling in real historical recordings due to unaddressed spectral properties, pushing the samples towards out-of-distribution regions compared to the training data and thus reducing the reliability of diffusion model predictions.
@@ 3.1 Filter Parameterization
The lowpass filter operator, initially applied in the BABE framework [7], is extended to a parametric frequency-response equalizer. The equalizer configuration, depicted in Fig. 1, incorporates an adjustable anchor cutoff frequency ($f_{0}$) maintained at 0 dB to ensure stability in the filter’s gain across its operational bandwidth. The design includes a set of adjustable slopes and breakpoints above this anchor frequency ($f_{i}$ and $A_{i}$), similar to those in BABE but with the flexibility of non-strictly decaying slopes. Beyond a certain point, marked by the last optimizable breakpoint and a fixed steep slope set to $A_{\text{lim+}}=-80$ dB, it is assumed that the band has no significant spectral content. This structure is symmetrically mirrored below the anchor cutoff, featuring adjustable slopes and breakpoints ($f_{-i}$ and $A_{-i}$), culminating in a fixed steep slope $A_{\text{lim-}}=80\,dB$ at the lowest frequency band to delineate the effective spectral range. The filter is piecewise differentiable with respect to the cutoff frequencies and slope parameters.
Choosing the right number of filter stages is crucial, especially under the assumption that the magnitude response should exhibit smoothness. Employing a filter with more stages provides greater flexibility during optimization but risks overfitting to minor spectral details. This overfitting can lead to an unrealistic magnitude response that could negatively impact the optimization. With this in mind, a smaller number of stages is favored to maintain smoothness, and we have chosen a configuration with 5 breakpoints.
@@
To mitigate potential training instabilities, we impose restrictions on the slopes, ensuring that they do not exceed predefined minimum and maximum values of -40 dB and 40 dB, respectively. Additionally, the cutoff frequencies are confined within the range $(f_{\text{min}},f_{\text{max}})$, which is determined by the FFT size employed in the filter parameterization. Furthermore, we enforce that a sequential ordering of the cutoff frequencies, requiring each frequency to be less than the subsequent one. These constraints are implemented through the application of projections within the specified ranges during the optimization process. Ideally, the constraint on the ordering of cutoff frequencies would be unnecessary, as the regularization scheme introduced in the next subsection aims to prevent parameter configurations that could lead to such scenarios.
@@ 3.4 Long-term Average Spectrum-based Initialization
A known blind equalization technique designs a filter to align the input LTAS with that off a reference set [4]. To achieve this, we calculate the reference LTAS, denoted as $\tilde{\mathbf{R}}_{\text{LTAS}}$, by applying the Short-Time Fourier Transform (STFT) to the reference set and average the power spectra across frames. In the STFT, we use 4096-sample windows at the sampling rate of $f_{s}=44.1$ kHz, or 2048-sample windows at $f_{s}=22.05$ kHz. We employ a hop size of 1/4 in both cases. One-third-octave band Gaussian smoothing is applied to the averaged spectra. The LTAS of the input $\tilde{\mathbf{Y}}_{\text{LTAS}}$ is determined in the same way, based on the original recording.
@@
Considering these limitations, we adopt this method as a preliminary step that BABE-2 will refine. Our goal is to facilitate the equalization process, aiming for a more stable optimization of the filter at the initial stages and achieving convergence with fewer iterations. We explore two approaches to accomplish this. Firstly, we utilize $H_{\text{LTAS}}$ for an improved initial setup of the audio signal at starting time $T$, $\mathbf{x}_{T}$. Unlike in BABE [7], where $\mathbf{x}_{T}$ began as a noisy version of the observations $\mathbf{x}_{T}\sim\mathcal{N}({\mathbf{y}},\sigma\text{start}^{2}\mathbf{I})$, we now introduce a modified approach. We suggest beginning with a noisy version of the equalized observations $\mathbf{x}_{T}\sim\mathcal{N}({\tilde{\mathbf{y}}},\sigma\text{start}^{2}\mathbf{I})$, where $\tilde{\mathbf{y}}$ is derived using Eq. (10). This version of our method is hereon referred to as “BABE-2 w/ LTAS-EQ init.”
The second strategy extends the use of LTAS-EQ beyond initial setup by incorporating it into the optimization objective. This involves replacing the original recording $\mathbf{y}$ with $\tilde{\mathbf{y}}$ in the reconstruction cost functions from Eqs. (12) and (LABEL:eq:filter_rec). It is crucial to recognize that in this context, the estimated frequency response $H_{\phi}$ no longer corresponds directly to the original recording response $\mathbf{y}$. Instead, the connection between the original recording and the estimated restored version $\hat{\mathbf{x}}_{0}$ becomes a composite of $H\phi$ and $H_{\text{LTAS}}$ as formalized below:
@@ 3.5 Inference Algorithm
$\tilde{\mathbf{y}}=\mathcal{F}^{-1}(H_{\text{LTAS}}^{-1}\odot\mathcal{F}(\mathbf{y})$) $\triangleright$ LTAS-based init. (Sec. 3.4)
if use LTAS-EQ in objective then
     $\mathbf{y}\leftarrow\tilde{\mathbf{y}}$
end if
Sample $\mathbf{x}_{T}\sim\mathcal{N}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\tilde{\mathbf{y}}},\sigma_{\text{start}}^{2}\mathbf{I})$triangleright apply warm initialization
@@
     $\phi_{i}^{0}\leftarrow\phi_{i+1}$triangleright use the filter from last step
     for $j\leftarrow 0,\dots,M_{\text{max its.}}$ do $\triangleright$ filter optimization
         $\hat{\mathbf{y}}_{\phi_{i}^{j}}\leftarrow\mathcal{F}^{-1}(H_{\phi_{i}^{j}}\odot\mathcal{F}(\mathbf{\hat{x}}_{0}))$triangleright apply filter
         $\mathcal{C_{\text{filter}}}(\phi_{i}^{j})\leftarrow\mathcal{C_{\text{rec}}}(H_{\phi_{i}^{j}}(\hat{\mathbf{x}}_{0}),\mathbf{y}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+\gamma_{\text{BCR}}\mathcal{C}_{\text{BCR}}(\phi_{i}^{j})}$
         $\phi_{i}^{j+1}\leftarrow\phi_{i}^{j}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathrm{AdamGrad}}(C_{\text{filter}}(\phi_{i}^{j}))$triangleright optim. step
         $\phi_{i}^{j+1}\leftarrow\mathrm{project}(\phi_{i}^{j+1})$triangleright project the filter params.
     end for
     $\phi_{i}\leftarrow\phi_{i}^{M}$
     $\hat{\mathbf{y}}_{\phi_{i}}\leftarrow\mathcal{F}^{-1}(H_{\phi_{i}}\odot\mathcal{F}(\mathbf{\hat{x}}_{0}))$triangleright apply filter
     $\bar{\mathbf{y}}=\mathbf{y}+\gamma\bm{\epsilon}$triangleright noise regularization (Sec. 3.3)
     $\mathbf{g}_{i}\leftarrow-\xi(\sigma_{i})\;\nabla_{\mathbf{x}_{i}}C_{\text{audio}}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bar{\mathbf{y}}},\mathbf{\hat{y}}_{\phi_{i}})$triangleright rec. guidance
     $\mathbf{s}_{i}\leftarrow\frac{\mathbf{\hat{x}}_{0}-\mathbf{x}_{i}}{\sigma_{i}^{2}}$triangleright prior score
     $\mathbf{x}_{{i-1}}\leftarrow\mathbf{x}_{i}-\sigma_{i}(\sigma_{i-1}-\sigma_{i})(\mathbf{s}_{i}+\mathbf{g}_{i})$triangleright update step
@@
The inference algorithm of BABE-2 shares strong similarities with the one employed in BABE [7]. The pseudocode in Algorithm 1 summarizes the inference process, highlighting the the new contributions in blue. Despite, for brevity reasons, Algorithm 1 shows an adaptation of a $1^{\text{st}}$ order Euler sampler, we use the $2^{\text{nd}}$ order stochastic sampler introduced in [9]. One relevant addition is the use of the Adam optimizer [14] to optimize the filter parameters, instead of a simpler stochastic gradient descent as used in [7]. Adam applies adaptive per-parameter learning rates and second-order moments of the gradients, which stabilizes the performance for non-convex optimization problems with sparse gradients.
@@
We also use the same reconstruction cost function as BABE for optimizing the filter parameters, which consists of a frequency-weighted $L_{2}$ norm between spectral magnitudes, denoted as:
$C_{\text{filter}}(\mathbf{y},H_{\phi}(\hat{\mathbf{x}}_{0}))=\left\lVert\mathbf{W}(|\mathcal{F}(\mathbf{y})|-|\mathcal{F}(H_{\phi}(\hat{\mathbf{x}}_{0}))|)\right\rVert_{2}^{2},$ (13)
where the matrix $\mathbf{W}$ denotes a frequency-dependent weighting function. Further implementation details are omitted from the main paper, but they are given in the Appendix A  [^2] The appendix is available in the pre-print: (link to arxiv here) and in the source code 333github.com/eloimoliner/BABE2.

### v2

@@ Abstract
This paper presents a novel approach to audio restoration, focusing on the enhancement of low-quality music recordings, and in particular historical ones. Building upon a previous algorithm called BABE, or Blind Audio Bandwidth Extension, we introduce BABE-2, which presents a series of improvements. This research broadens the concept of bandwidth extension to generative equalization, a novel task that, to the best of our knowledge, has not been explicitly addressed in previous studies. BABE-2 is built around an optimization algorithm utilizing priors from diffusion models, which are trained or fine-tuned using a curated set of high-quality music tracks. The algorithm simultaneously performs two critical tasks: estimation of the filter degradation magnitude response and hallucination of the restored audio. The proposed method is objectively evaluated on historical piano recordings, showing an enhancement over the prior version. The method yields similarly impressive results in rejuvenating the works of renowned vocalists Enrico Caruso and Nellie Melba. This research represents an advancement in the practical restoration of historical music. Historical music restoration examples are available at: research.spa.aalto.fi/publications/papers/dafx-babe2/.
@@ 1 Introduction
Historical music recordings suffer from severe impairment due to limitations of the physical recording media, alongside the wear and tear from playback and storage [1, 2]. These degradations can take the form of additive disturbances, decreased bandwidth, coloration, and nonlinear effects [3]. This project goal is to enhance the quality of historical music recordings the audio standards of today, making them more accessible for people used to modern sound quality. Restoring these recordings is challenging for several reasons. First, the problem is ill-posed, , with a multitude of potential solutions, thereby complicating the quest for the optimal one. Additionally, the degradation process is unknown, making the restoration a blind inverse problem for which the solution must be found without definite information on the necessary repairs.
In the 1970s, Stockham et al. began groundbreaking efforts in digitally restoring gramophone recordings, notably those of the celebrated singer Enrico Caruso [4]. At the time, the approach was limited by technological constraints, relying on the estimation of an equalization curve derived from the spectral average of a contemporary recording. Since then, the field of audio restoration has evolved significantly [1]. Today, data-driven techniques employing deep neural networks can remove all additive disturbances simultaneously [5]. Deep generative models, such as diffusion models [6], address complex restoration tasks, such as bandwidth extension [7], with unprecedented success. This study investigates a critical question: “Can deep generative models elevate the quality of historical music recordings to modern standards?
To aid in this investigation, we introduce and explore the concept of generative equalization, a task designed to overcome the limitations of traditional equalization on ill-conditioned scenarios, i.e., when some frequency bands are not present in the original recording. In such cases, the equalization curve must be cropped; otherwise, it will only amplify background noise. In contrast, a generative equalizer can synthesize the missing spectral components as to achieve the target spectral profile, as it is a generalization of bandwidth extension. While the combination of equalization and bandwidth extension has been previously studied [Qian2004combining], this paper, to the best of our knowledge, is the first to explore this approach in the context of generative models and music restoration.
@@
This paper is structured as follows. Sec. 2 outlines the basic principles required to follow the description of the new algorithm. This section describes the basics of diffusion models, diffusion posterior sampling, and the previous BABE algorithm. Following this, Sec. 3 delves into the enhancements made in the BABE-2 method, expanding the bandwidth extension techniques introduced in [7] to encompass the broader concept of generative equalization. Sec. 4 reports on experiments conducted with historical recordings of piano and singing voice, and Sec. 5 describes our procedure for selecting the training data used for restoring specific historical singers. Furthermore, this section elucidates on some of the capabilities and limitations of BABE-2 for singing voice restoration. Finally, Sec. 6 concludes the paper.
@@ 2.1 Diffusion Models
$d\mathbf{x}_{\tau}=-\tau\nabla_{\mathbf{x}_{\tau}}\log p_{\tau}(\mathbf{x}_{\tau})d\tau.$ (1)
@@
$\mathbb{E}_{\mathbf{x}_{0}\sim p_{\text{data}},\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\lambda(\tau)\lVert D_{\theta}(\mathbf{x}_{0}+\tau\mathbf{\epsilon},\tau)-\mathbf{x}_{0}\rVert_{2}^{2}\right],$ (3)
where $\lambda(\tau)$ is a weighting parameter defined according to [10]. In the subsequent sections, the term $\hat{\mathbf{x}}_{0}$ denotes the denoised estimate obtained through the application of neural network $D_{\theta}$.
@@ 2.2 Diffusion Posterior Sampling
where $\xi(\tau)$ serves as a gradient normalization factor [11], and $C_{\text{audio}}(\cdot,\cdot)$ denotes an appropriately chosen cost function.
Within this framework, $\hat{\mathbf{x}}_{0}=D_{\theta}(\mathbf{x}_{\tau},\tau)$ represents the Minimum Mean Squared Error (MMSE) estimate of the clean audio $\mathbf{x}_{0}$ given the noisy or transformed state $\mathbf{x}_{\tau}$. It is important to note that implementing this method necessitates differentiating through the denoiser, which is a deep neural network. The differentiation step adds a layer of computational complexity to the process. This approach has been successfully used for solving audio restoration tasks, such as bandwidth extension, inpainting, and declipping [11, hernandez2024vrdmg].
@@ 2.3 Application to Blind Inverse Problems
In BABE, the filter parameters are constrained to ensure they describe a strictly decaying function [7]. The reason for doing that was to limit the filter design to a lowpass characteristic, aiming to stabilize the optimization convergence. Such a parameterization, however, does not accommodate spectral coloration or resonances, which are common in historical recordings and can significantly impact the audio quality. We hypothesize that the constrained modeling capability degrades the performance of posterior sampling in real historical recordings due to unaddressed spectral properties, pushing the samples towards out-of-distribution regions compared to the training data and thus reducing the reliability of diffusion model predictions.
@@ 3.1 Filter Parameterization
The lowpass filter operator, initially applied in the BABE framework [7], is extended to a parametric frequency-response equalizer. The equalizer configuration, depicted in Fig. 1, incorporates an adjustable anchor cutoff frequency ($f_{0}$) with a fixed magnitude at 0 dB to ensure stability in the filter’s gain across its operational bandwidth. The design includes a set of adjustable breakpoints and slopes ($f_{i}$ and $A_{i}$) localized above and below the anchor frequency, similarly to those employed in BABE but with the flexibility of non-strictly decaying slopes. The first and last breakpoints in in Fig. 1 delineate the effective spectral range, assuming there is no significant spectral energy beyond these limits. The slopes before and after the limiting breakpoints are not optimizable and are fixed to $A_{\text{lim-}}=80$ dB and $A_{\text{lim+}}=-80$ dB, respectively. We refer to Appendix A.4  [^2] Appendix available in the pre-print: arxiv.org/abs/2403.18636 for further details on the filter paramaterization. The filter is piecewise differentiable with respect to the cutoff frequencies and slope parameters.
Choosing the right number of filter stages is crucial, especially under the assumption that the magnitude response should exhibit smoothness. Employing a filter with more stages provides greater flexibility during optimization but risks overfitting to minor spectral details. This overfitting can lead to an unrealistic magnitude response that could negatively impact the optimization. With this in mind, a smaller number of stages is favored to maintain smoothness, and we have chosen a configuration with 5 breakpoints, as depicted in Fig 1.
@@
To mitigate potential training instabilities, we impose restrictions on the slopes, ensuring that they do not exceed predefined minimum and maximum values of -40 dB and 40 dB, respectively. Additionally, the cutoff frequencies are confined within the range $(f_{\text{min}},f_{\text{max}})$, which is determined by the FFT size employed in the filter parameterization. In our experiments, we set $f_{\text{min}}=10$ Hz and $f_{\text{max}}=\frac{1}{2}f_{\text{s}}$ to the Nyquist limit, depending on the sampling rate. Furthermore, we enforce that a sequential ordering of the cutoff frequencies, requiring each frequency to be less than the subsequent one. These constraints are implemented through the application of projections within the specified ranges during the optimization process. Ideally, the constraint on the ordering of cutoff frequencies would be unnecessary, as the regularization scheme introduced in the next subsection aims to prevent parameter configurations that could lead to such scenarios.
@@ 3.4 Long-term Average Spectrum-based Initialization
A known blind equalization technique designs a filter to align the input LTAS with that of a reference set [4]. To achieve this, we calculate the reference LTAS, denoted as $\tilde{\mathbf{R}}_{\text{LTAS}}$, by applying the Short-Time Fourier Transform (STFT) to the reference set and average the power spectra across frames. In the STFT, we use 4096-sample windows at the sampling rate of $f_{s}=44.1$ kHz, or 2048-sample windows at $f_{s}=22.05$ kHz. We employ a hop size of 1/4 in both cases. One-third-octave band Gaussian smoothing is applied to the averaged spectra. The LTAS of the input $\tilde{\mathbf{Y}}_{\text{LTAS}}$ is determined in the same way, based on the original recording.
@@
Considering these limitations, we adopt this method as a preliminary step that BABE-2 will refine. Our goal is to facilitate the equalization process, aiming for a more stable optimization of the filter at the initial stages and achieving convergence with fewer iterations. We explore two approaches to accomplish this. Firstly, we utilize $H_{\text{LTAS}}$ for an improved initial setup of the audio signal at starting time $T$, $\mathbf{x}_{T}$. Unlike in BABE [7], where $\mathbf{x}_{T}$ began as a noisy version of the observations $\mathbf{x}_{T}\sim\mathcal{N}({\mathbf{y}},\sigma_{\text{start}}^{2}\mathbf{I})$, we now introduce a modified approach. We suggest beginning with a noisy version of the equalized observations $\mathbf{x}_{T}\sim\mathcal{N}({\tilde{\mathbf{y}}},\sigma_{\text{start}}^{2}\mathbf{I})$, where $\tilde{\mathbf{y}}$ is derived using Eq. (10). This version of our method is hereon referred to as “BABE-2 w/ LTAS-EQ init.”
The second strategy extends the use of LTAS-EQ beyond initial setup by incorporating it into the optimization objective. This involves replacing the original recording $\mathbf{y}$ with $\tilde{\mathbf{y}}$ in the reconstruction cost functions from Eqs. (12) and (13). It is crucial to recognize that in this context, the estimated frequency response $H_{\phi}$ no longer corresponds directly to the original recording response $\mathbf{y}$. Instead, the connection between the original recording and the estimated restored version $\hat{\mathbf{x}}_{0}$ becomes a composite of $H\phi$ and $H_{\text{LTAS}}$ as formalized below:
@@ 3.5 Improved Inference Algorithm
$\mathbf{x}_{T}\leftarrow$ Warm Initialization$(\mathbf{y})$triangleright Initialize audio signal
@@
     Optimize $\phi_{i}$triangleright Optimization loop
     Likelihood score approximation ($\approx\nabla_{\mathbf{x}_{i}}\log p(\mathbf{y}|\mathbf{x}_{i},\phi_{i})$)
     Update $\mathbf{x}_{i}$
@@
The inference algorithm of BABE-2 shares strong similarities with the one employed in BABE [7]. The pseudocode in Algorithm A summarizes the inference process, highlighting the new contributions in blue. Despite, for brevity reasons, Algorithm 1 shows an adaptation of a $1^{\text{st}}$ order Euler sampler, we use the $2^{\text{nd}}$ order stochastic sampler introduced in [9]. One relevant addition is the use of the Adam optimizer [15] to optimize the filter parameters, instead of a simpler stochastic gradient descent as used in [7]. Adam applies adaptive per-parameter learning rates and second-order moments of the gradients, which stabilizes the performance for non-convex optimization problems with sparse gradients.
@@
In contrast, we now employ a slightly different reconstruction cost function for optimizing the filter parameters. BABE-2 uses a high-frequency emphasized $L_{2}$ norm, denoted as
$C_{\text{filter}}(\mathbf{y},H_{\phi}(\hat{\mathbf{x}}_{0}))=\lVert H_{\text{emph.}}(\mathbf{y}-H_{\phi}(\hat{\mathbf{x}}_{0}))\rVert_{2}^{2},$ (13)
where $H_{\text{emph.}}$ is a first-order high-pass pre-emphasis filter [16]. Further implementation details are omitted from the main paper, but they are given in the Appendix A and in the source code 333github.com/eloimoliner/BABE2.